1. This process is also known as combustion. In a laboratory, a 0.5484 g glucose sample underwent combustion, producing carbon dioxide gas and liquid water.

a. Write the balanced chemical equation for the reaction described above.

b. How many moles of glucose are present before the reaction?

c. What is the mole relationship between glucose and carbon dioxide? Between glucose and water?

d.  How many grams of CO2 is produced from the given amount of sample.







1
Expert's answer
2021-10-28T03:01:52-0400

a) "C_6H_{12}O_6+6O_2\to 6CO_2+6H_2O"

b) RMM for glucose ="12\times 6+1\times 12+(16\times6)=180"

moles of glucose"=\frac{0.5484}{180}=0.003047moles"

c) mole ratio "C_6H_{12} : 6CO_2 =1:6"

mole ratio "C_6H_{12}:6H_2O=1:6"

d) mole of "CO_2=0.003047\times 6=0.018282moles"

"RMM \space of \space CO_2=12+16\times 2=48"

mass of "CO_2=48\times 0.018282=0.8775g"